OpenStudy (annie96):

1. Solve the equation? 2(5m+3)-8m=4 2. Write a equation to represent the given facts below. A number times 6 decreased by 4 is -40. solve the problem above. show all work. How do you check your answer to the equation above? Explain.

OpenStudy (anonymous):

2(5m+3)-8m=4 10m +6-8m=4 2m=-2 m=-1

OpenStudy (anonymous):

A number times 6 decreased by 4 is -40 let x be the number so 6x-4= -40 x= -6